# Network Formation and Dynamics Among Multi-LLMs

### v2

@@ Abstract.
Social networks shape opinions, behaviors, and information dissemination in human societies. As large language models (LLMs) increasingly integrate into social and professional environments, understanding their behavior within the context of social interactions and networks becomes essential. Our study analyzes LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics. We observe that LLMs exhibit key social network principles, including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon, when asked about their preferences in network formation. We also investigate LLMs’ decision-making based on real-world networks, revealing that triadic closure and homophily have a stronger influence than preferential attachment and that LLMs perform well in network formation predictions. Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their network formation behaviors and exploring their impacts on social dynamics.
@@ 1. Introduction
Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023) and LLaMA2 (Touvron et al., 2023), have shown promising developments in AI techniques and their integration into real-life applications. It is thus crucial to comprehend AI actions to ensure they align with human expectations, mitigate potential risks, and maximize their benefits. Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019). Recently, researchers have started to apply methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs. These methods help uncover the generative capabilities of LLMs and their interpretability (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
@@
In human societies, social networks play a crucial role in shaping individual behaviors, preferences, and connections, as well as influencing the diffusion of information and norms across communities (Rogers, 2003; Bakshy et al., 2012; Fowler and Christakis, 2008; Banerjee et al., 2013; Yuan et al., 2018). LLMs have shown great potential in social contexts, notably as intelligent personal assistants facilitating social and business interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)). However, less is known about how LLMs’ innate behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023). This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the intricate networks of human relationships, which is a fundamental aspect of social systems.
@@
Our study thus explores LLMs’ behaviors and preferences in network formation in both synthetic and real-world social networks. By analyzing interactions between multiple LLMs (or multi-LLMs), we aim to understand the implications of LLMs representing humans in social and business settings. Specifically, we examine micro-level social network properties including preferential attachment (Barabási and Albert, 1999), triadic closure (Granovetter, 1973), and homophily (McPherson et al., 2001), as well as macro-level properties including community structure (Newman and Girvan, 2004), and the small-world phenomenon (Kleinberg, 2000; Watts and Strogatz, 1998).
@@
In synthetic networks, our findings reveal that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions. Moreover, LLMs demonstrate strong tendencies toward triadic closure and exhibit tendency of homophily. These tendencies further explain the emergence of community structures with high modularity in networks formed by multiple LLMs. Additionally, we show that LLM agents exhibit dynamics consistent with small-world phenomena akin to those observed in the Watts-Strogatz model (Watts and Strogatz, 1998).
@@
Next, we conduct a comparative analysis of these principles to identify the key factors driving network dynamics in the context of real-world network formation scenarios. To do so, we consider real-world networks and ask each LLM to simulate a node in the network. We provide information on their network structure and ask for their decisions for the next link formation or dissolution step. We approach this problem as a discrete choice modeling problem (Liben-Nowell and Kleinberg, 2003; Overgoor et al., 2019). In this real-world setting, we find that LLMs’ choices are primarily weighted on homophily and triadic closure, compared to preferential attachment, and that LLMs substantially exceed random guessing when predicting new links. Additionally, the resulting choices of LLM agents strengthen the network’s community structure.
@@
Our study contributes to the literature on LLMs in several ways. First, it advances the discussion on aligning LLMs with human social preferences, which has potential applications in developing socially aware LLMs for real-life interactions. Insights into LLMs’ innate behaviors in network formation can inform the design of LLMs for effective collaboration in social and professional settings. Second, our work has direct applications to synthetic dataset generation, a problem important for the graph learning community. By controlling for a few parameters, such as the temperature, our work can be utilized to generate a variety of datasets, as well as generate synthetic datasets from real-world data to protect privacy. Finally, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings.
ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$ formatted as a Python f-string.
@@
Figure 1. Networks with $n=200$ nodes created based on Principle 1 for various temperatures. For the structure-based networks (Figure 1(a)), we provide the multi-LLMs with the neighborhoods $\{N_{j,t}:j\in i_{t}\}$, and for the degree-based networks (Figure 1(b)), we provide the degrees $\{d_{j,t}:j\in i_{t}\}$ of the nodes $j\in i_{t}$. We also plot the degree distribution and compare it against a BA network with $n=200$ nodes. Top: In Figure 1(a), we observe that the resulting networks follow a power-law degree distribution ($P>0.5$, Bootstrap hypothesis test). When the temperature is 1.5, the resulting network resembles the BA model ($P<0.001$, K-S test). Bottom: When the degree information is provided only (Figure 1(b)), the resulting networks correspond to unrealistic star-like formations that are far from a scale-free network and resemble more a core-periphery network.
2. Results
Our study employs the state-of-the-art LLMs, such as GPT-3.5 and GPT-4 (OpenAI, 2023) with the OpenAI API [^1] For GPT-3.5 we use the gpt-3.5-turbo model provided by OpenAI, and for GPT-4 we use the gpt-4-1106-preview model.. We constructed multiple LLMs using multiple independent conservation threads where each thread represents one multi-LLM. We prompt each LLM by providing information about other LLMs’ network information as well as decisions they are faced with. We test both state-of-the-art GPT-based models: for the results of Sections 2.2 and 2.1, we use GPT-3.5 agents, and in Section 2.3, we use GPT-4 agents. We obtained similar results using GPT-3.5 and GPT-4. Overall, we adopted a standard framework for multi-LLMs but underscored the multi-agent nature of the design in our setting (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023). We evaluate the robustness of our results by testing various temperature settings in LLMs, which control the degree of variation of the model output. All experiments employ a dynamic network setup. We also employ an interview-like approach to inquire about the rationale behind the decisions of LLMs. The details of our methods are presented in Methods.
In this section, we first discuss three major micro-level network principles: preferential attachment, triadic closure, and homophily. We then examine two macro-level network principles: community structure and small-world phenomenon. Subsequently, we test the strength of these principles when LLM agents face real-world networks and compare the preferences in network formation tasks between LLMs and humans.
2.1. Micro-Level Properties
@@
Preferential attachment is a well-known mechanism in network science that explains how certain nodes in a network become more connected over time, leading to the emergence of highly connected vertices and scale-free distribution of degree (Barabási and Albert, 1999; Bianconi and Barabási, 2001).
We examine whether LLMs exhibit the preferential attachment mechanism in the following way: First, we start with an empty network and add nodes at each time step $t$. For each new time step, we add a new node $i_{t}$, providing information on the network neighbors of all existing nodes. [^2] See Appendix E, for information about how the features are represented in the prompt. We chose to generate networks with $n=200$ nodes, as this size is sufficiently large to allow for the observation of degree distribution patterns.
We plot one of the examples of resulting networks under each temperature in Figure 1(a). As observed in all temperature setups, we observe a few nodes that have many connections, and the majority of the notes have one or few connections, which renders the degree distribution scale-free, i.e., the probability $\pi(d)$ a node has degree $d$ is given by
$\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
Furthermore, we observe mode “super nodes” with many connections as the temperature increases.
We also compare with the Barabási–Albert (BA) model (Barabási and Albert, 1999) where each incoming node $t$, the probability $p_{i,t}$ of it connecting to node $i$ is proportional to ${d_{i,t}}$, where $d_{i,t}$ is the degree of node $i$ at time step $t$, namely,
$p_{i,t}=\frac{d_{i,t}}{\sum_{j=1}^{t}d_{j,t}}.$
For temperatures of 0.5 and 1.0, we find that the estimated exponent values are $\hat{\gamma}=1.50\pm 0.23$ and $1.92\pm 0.24$, respectively. However, the degree distributions do not resemble the degree distribution of the generated BA network with an expected exponent $\gamma=3$. This is supported by the $P$-values from a two-sample Kolmogorov-Smirnov (K-S) test, which are $P<0.001$ and $P<0.05$, respectively when compared to the empirical degree distribution of the BA network [^3] As shown by (Barabási, 2013), the BA model follows a scale-free distribution of the form $\pi_{\mathrm{BA}}(d)\propto d^{-3}$ for sufficiently large $n$. However, because our $n$ is small, so that the prompt does not exceed the token limit for the GPT-3.5 model, we directly compared it to a sample from the BA model (the generated network has $\hat{\gamma}_{\mathrm{BA}}=2.24\pm 0.24$), instead of comparing with the limiting distribution $\pi_{\mathrm{BA}}(d)$..
Moreover, for all temperatures, the resulting degree distribution is scale-free ($P>0.5$, bootstrapping hypothesis test comparing with the fitted scale-free distribution).
We also examine the network formation when only degree information (rather than network structure) is provided, in a similar spirit of (De Marzo et al., 2023). We present the results in Figure 1(b). We observe “star-like” formations around one node. The degree distributions differ from scale-free distributions produced by statistical models such as the BA model ($P<0.001$, two-sample K-S test compared to the empirical distribution of the generated BA network). They are closer to a core-periphery network (Borgatti and Everett, 2000).
Finally, we employ text analysis of reasoning provided by the agents for the results of Figure 1(a) [^4] We do not provide the distribution of reasons for the results of Figure 1(b) as all of the LLM’s decisions are made given the degree information.. We observe two main categories of reasons (see Appendix C) for the fraction under each temperature):
  • Well-Connectedness: it corresponds to reasons related to the person being well-connected and having many friends.
  • Mutual Friends: it is related to having many friends in common with the person so that the person can introduce them to their friends.
The well-connectedness reason becomes less prevalent as the temperature increases, and the mutual friends’ reason becomes more prevalent. Note that when the temperature is 1.5, the two reasons have almost the same frequency.
Taken together, networks created by LLM systems, when given a network structure, demonstrate preferential attachment and exhibit scale-free degree distributions.
@@
Figure 2. Networks created based on Principle 2. Top: Figure 2(a) shows the resulting networks created by Principle 2 when the intersection of the neighborhoods of the query node and each alternative is provided. The node colors correspond to the stochastic block model (SBM) groups each node belongs to. The red edges correspond to the newly created edges between nodes of the red cluster, the blue edges correspond to the newly created edges between nodes of the blue cluster, and the orange edges correspond to edges created between the two clusters. The newly formed edges are drawn with a larger line width. We observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). Bottom: Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative. Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test).
@@
The second micro-level principle we examine is triadic closure, which describes the phenomenon that individuals are likely to form connections with their friends’ friends, leading to closed triads within the network. This tendency enhances the network’s structure and cohesion, rooted in the likelihood of two nodes forming a direct connection if they share a common neighbor (Granovetter, 1973; Mosleh et al., 2024).
@@
To test for triadic closure, we use an assortative stochastic block model (SBM) (Newman, 2003) to generate an initial seed network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$. Connections within each cluster have a probability of $0.5$, while connections across clusters have a probability of $0.1$. This setup reflects our intuition that nodes within the same cluster are more likely to form connections due to their higher number of common neighbors. We then iterate over each node $i$, providing features based on the intersection of the neighborhoods of all of $i$’s non-neighbors or the number of common friends between $i$ and any non-neighbor $j$. This modeling choice contrasts the results mentioned about preferential attachment, where the neighborhood information is provided and made to make the two settings consistent [^5] We want to note that performing the same experiment when we provide the neighbors instead of the common neighbors produces similar results..
We compare this with the random null model, which selects a node at random to connect to. We run ten simulations for $n=50$ nodes. [^6] For better visualization, we chose $n=50$ nodes instead of a large number of nodes such as $n=200$. Moreover, the results are statistically significant even with $n=50$ nodes.
To measure whether the principle of triadic closure exists, we employ the fraction $\hat{p}$ of newly added edges between nodes of the same community, i.e., the fraction of edges in $G_{T}\setminus G_{1}$ that have their ends both at either $A$ or $B$, namely
@@
where $y_{i},y_{j}\in\{A,B\}$ are the community memberships of nodes $i$ and $j$ respectively, and $E(G_{T})\setminus E(G_{1})$ corresponds to the newly added edges.
Generally, a value of $\hat{p}>0.5$ indicates the tendency of triadic closure since an LLM agent, which is supposed to exhibit triadic closure given the common neighbors, would ideally create a link within the same cluster.
Additionally, we introduce the concept of marginal transitivity for the network, represented by $D$. This metric is defined as the change in the proportion of closed triangles to open triads in the network, from $G_{T}$ (the initial network) to $G_{1}$ (which corresponds to the SBM):
We depict the sample outcomes in Figure 2(a), where the upper panel presents the results when the entire network structure is provided, whereas the lower panel only provides the number of common neighbors. Node colors correspond to their respective groups within the SBM. Red edges denote connections within the red cluster, blue edges within the blue cluster, and orange edges between the two clusters. Newly established edges are emphasized with a thicker line width.
We present the probabilities of network formation within a community in the last panels in each row. We find that the probability of forming a within-community tie across all temperatures is substantially larger than the random network formation scenario ($P<0.001$, t-test comparing to Random), suggesting LLMs’ tendency of triadic closure (since an LLM which is given the common neighbors as features should ideally create links within the same community).
Additionally, we observe a statistically significant increase in the marginal transitivity $D$ for all temperatures ($P<0.001$, t-test). Moreover, in all cases, the marginal transitivity $D$ is statistically significant compared to Random ($P<0.001$, t-test).
These results suggest that LLM agents exhibit the tendency of triadic closure. Similar to the approach in earlier reasoning analysis, we present the most common reasons provided by LLMs in Appendix C [^7] We do not provide the distribution of reasons for the results of Figure 2(b) as all of the LLM’s decisions are made given the number of common neighbors information.. We find that most of the agents base their decisions on the number of mutual friends.
@@
(a) Assortativity Coefficients
(b) Network Instances
(c) Community Structure
Figure 3. Top: Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily. The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random) Middle: Networks created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. The edge widths correspond to the number of common attributes shared between the nodes. We report the probability of forming a link as a function of the number of common attributes between two nodes. The node colors correspond to the communities found by running the Louvain algorithm. Bottom: Louvain modularity, number of communities, and community sizes of networks created based on Principle 3. Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test). For smaller temperatures, the communities are more “well-separated”.
@@
To measure the presence of homophily, we measure the attribute assortativity coefficient of each of the three features (hobby, color, location). For a property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$, its assortativity coefficient $R$ is defined as
$R=\frac{\sum_{i=1}^{n}M_{ii}-\sum_{i=1}^{n}a_{i}b_{i}}{1-\sum_{i=1}^{n}a_{i}b_{i}},$
where $M$ is the mixing matrix with entries $M_{ij}$ corresponding to the fraction of edges where one end has value $P_{i}$, and the other end has value $P_{j}$, $a_{i}=\sum_{j=1}^{n}M_{ij}$ and $b_{i}=\sum_{j=1}^{n}M_{ji}$. The assortativity has values between $-1$ and $+1$, where a positive value denotes the tendency of the nodes to connect to other similar nodes (hence yielding a homophilous network), and a negative value corresponds to connecting to the tendency to connect to dissimilar nodes (which indicates heterophily).
In Figure 3, we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures. Moreover, the observed assortativity at all temperatures significantly deviates from Random, which has assortativity close to 0 ($P<0.001$, t-test comparing with Random). Taken together, we conclude homophily emerges for the networks generated by LLMs.
Moreover, we present the probability of each reason being mentioned in Appendix C. We find that shared hobbies appear to be much more important than shared location or favorite color. This is consistent with our intuition that, similarly to humans, LLMs care more about hobbies than other superficial attributes when forming social connections.
2.2. Macro-Level Principles
@@
Both triadic closure and homophily contribute to the formation of community structures.By examining how these two factors contribute to network formation, we aim to gain insights into the underlying mechanisms driving community dynamics in LLM-generated networks. We employ the simulation results presented in Section 2.1 to determine whether community structure in networks generated by LLMs emerges from triadic closure or homophily.

### v3

@@ Abstract.
Social networks shape opinions, behaviors, and information dissemination in human societies. As large language models (LLMs) increasingly integrate into social and professional environments, understanding their behavior within the context of social interactions and networks becomes essential. Our study analyzes LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics. We observe that LLMs exhibit key social network principles, including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon, when asked about their preferences in network formation. We also investigate LLMs’ decision-making based on real-world networks, revealing that triadic closure and homophily have a stronger influence than preferential attachment and that LLMs perform well in network formation predictions. Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their social interaction behaviors and exploring their impacts on social dynamics.
@@ Introduction
Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023) and Llama 2 (Touvron et al., 2023), have shown promising developments in AI techniques and their integration into real-life applications. It is thus crucial to comprehend AI actions to ensure they align with human expectations, mitigate potential risks, and maximize their benefits. Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019). Recently, researchers have started to apply social science methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs. These methods not only reveal the capabilities and interpretability of LLMs but also suggest their potential for applications in social science (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
@@
In human societies, social networks play a crucial role in shaping individual behaviors, preferences, and connections, as well as influencing the diffusion of information and norms across communities (Rogers, 2003; Bakshy et al., 2012; Fowler and Christakis, 2008; Banerjee et al., 2013; Yuan et al., 2018). LLMs have shown great potential in social contexts, notably as intelligent personal assistants that facilitate social and prosocial interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)). However, less is known about how LLMs’ innate behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023). This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the networks of human relationships, which is a fundamental aspect of social systems.
@@
Our study thus explores LLMs’ behaviors and preferences in the context of network formation with both synthetic and real-world social networks. By analyzing interactions between multiple LLMs (or multi-LLMs), we aim to understand the implications of LLMs representing humans in social and professional settings. Specifically, we examine micro-level social network properties including preferential attachment (Barabási and Albert, 1999), triadic closure (Granovetter, 1973), and homophily (McPherson et al., 2001), as well as macro-level properties including community structure (Newman and Girvan, 2004), and the small-world phenomenon (Kleinberg, 2000; Watts and Strogatz, 1998).
@@
In simulations conducted using synthetic networks, we have the following findings: First, we observe that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions. Moreover, LLMs demonstrate strong tendencies toward triadic closure and exhibit tendency of homophily. These tendencies further explain the emergence of community structures in networks formed by multi-LLMs. Additionally, we show that LLM agents exhibit dynamics consistent with small-world phenomena akin to those observed in the Watts-Strogatz model (Watts and Strogatz, 1998).
@@
Next, we compare between these principles to identify the key factors driving network dynamics in real-world network formation contexts. To achieve this, we consider real-world networks and ask each LLM to simulate a node in the network. We provide information on their network structure and ask for their decisions for the next link formation or dissolution step. We approach this problem as a discrete choice modeling problem (Liben-Nowell and Kleinberg, 2003; Overgoor et al., 2019). In this real-world setting, we find that LLMs’ choices are primarily weighted on homophily and triadic closure, compared to preferential attachment, and that LLMs substantially exceed random guessing when predicting new links. Additionally, the resulting choices of LLM agents strengthen the network’s community structure.
@@
Our study contributes to the literature on LLMs in multiple ways. First, it advances the discussion on aligning LLMs with human preferences, which has potential applications in developing socially aware LLMs for real-life interactions. Insights into LLMs’ innate behaviors in network formation can inform the design of LLMs for effective collaboration in social and professional settings. Second, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings. Finally, our work has direct applications to synthetic dataset generation, a problem important for the network science community. By controlling for a few parameters, such as the temperature, our work can be utilized to generate a variety of datasets, as well as generate synthetic datasets from real-world data to protect privacy.
@@
Figure 1. Results for Principle 1 (preferential attachment): We display simulated networks comprising 200 nodes across different temperatures. For structure-based simulations (Figure 1(a)), the multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$, while for degree-based simulations (Figure 1(b)), node degree data $\{d_{j,t}:j\in V_{t}\}$ was provided ($V_{t}$ corresponds to the vertex set of the network $G_{t}$ at round $t$). The degree distributions were then compared with those of a BA network with $n=200$ nodes. (a): Power-law degree distributions are evident ($P>0.5$, K-S test), with the networks at a temperature of 1.5 closely resembling the BA model ($P>0.1$, K-S test). (b): With degree information only, the networks form more unrealistic star-like structures, diverging from scale-free configurations and more closely mirroring a core-periphery network structure.
Results
In our study, we leveraged state-of-the-art LLMs through the OpenAI Application Programming Interface (API). Our methodology involved constructing multiple LLMs, each acting as an independent agent within separate conversational threads. Each LLM agent was prompted with network information about its peers and faced with decision-making scenarios. For synthetic network simulations, we employed GPT-3.5 agents, whereas GPT-4 agents were used for real-world network analysis. [^1] We utilized the gpt-3.5-turbo model for GPT-3.5 and the gpt-4-1106-preview model for GPT-4, both provided by OpenAI. Notably, both GPT-3.5 and GPT-4 yielded consistent results.
We adopt a multi-LLM framework, emphasizing the multi-agent aspect of our experimental design, which shares similar spirits with recent research in the field (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023). To assess the robustness of our findings, we varied the ‘temperature’ settings in LLMs, which modulate the output variability, in a dynamic network environment. Additionally, we adopted an interview-like method to probe the rationale behind the LLMs’ decision-making processes, enriching our understanding of their behavior in network simulations.
In our analysis, we begin with an examination of three fundamental micro-level network principles: preferential attachment, triadic closure, and homophily. Subsequently, we shift our focus to two macro-level network phenomena: community structure and the small-world phenomenon. Finally, we assess the resilience of these principles as LLM agents interact with real-world networks. This allows us to compare and contrast the network formation preferences between LLMs and humans.
Micro-Level Properties
@@ Results
Preferential attachment is a fundamental concept in network science, illustrating how nodes in a network gain connections over time, leading to a scale-free degree distribution characterized by a few highly connected nodes (Barabási and Albert, 1999; Bianconi and Barabási, 2001).
In our study, we investigate whether LLMs manifest the preferential attachment mechanism. Starting with an empty network, we introduce nodes sequentially at each time step $t$. A new node $i_{t}$ is added during time step $t$, with the network’s existing nodes providing context about their connections. We generated networks with $n=200$ nodes, a size large enough to observe meaningful patterns in the degree distribution. Our simulations include two distinct types: structure-based, where the network’s structure was provided, and degree-based, where only the degree information of potential connections was given. [^2] For details on how features were represented in the prompts, refer to SI Section 5.
The graphical outputs from our structure-based simulation, as illustrated in Figure 1(a), reveal a distinct pattern where a small number of nodes possess a high number of connections, whereas the majority exhibit fewer links. This pattern suggests a scale-free degree distribution, characterized by the probability $\pi(d)$ that a node has degree $d$, mathematically represented as:
(1) $\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
Interestingly, we also observe that the number of highly connected “super nodes” increases with temperature.
[…]
Furthermore, in degree-based simulations in a similar spirit of ((De Marzo et al., 2023); see Figure 1(b)), we noted formations resembling a “star” around central nodes. These distributions deviate from the scale-free patterns seen in statistical models like the BA model ($P<0.001$, two-sample K-S test), aligning more closely with core-periphery structures (Borgatti and Everett, 2000).
To interpret the LLM agents’ decision-making, we analyzed the reasoning they provided, focusing on the structure-based results (Figure 1(a)) [^4] The degree-based results’ reasoning is not included as decisions were based solely on degree information.. Two primary reasons emerged, varying with temperature (detailed in SI Section 3):
  • Well-Connectedness: Relating to the node’s degree (number of connections).
  • Mutual Friends: Relating to the number of shared connections.
Interestingly, as the temperature increased, the prevalence of the well-connectedness rationale decreased, while the mutual friends’ rationale increased.
In conclusion, our analysis demonstrates that LLMs, when provided with network structure, can replicate the preferential attachment phenomenon.
@@
Figure 2. Results for Principle 2 (triadic closure). (a) Figure 2(a) shows the resulting networks created by triadic closure when the intersection of the neighborhoods of the query node and each alternative is provided. The node colors correspond to the stochastic block model (SBM) groups to which each node belongs. The red edges correspond to the newly created edges between nodes of the red cluster, the blue edges correspond to the newly created edges between nodes of the blue cluster, and the orange edges correspond to edges created between the two clusters. The newly formed edges are drawn with a larger line width. We observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). (b) Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative. Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). The error bars correspond to 95% confidence intervals.
@@
The second micro-level principle we examine is triadic closure, which posits that individuals are more likely to form connections with friends of friends, thus creating closed triads in the network. This process strengthens network structure and cohesion, grounded in the idea that two nodes are more likely to connect if they share a common neighbor (Granovetter, 1973; Mosleh et al., 2024).
@@
To investigate triadic closure, we employ an assortative stochastic block model (SBM) (Newman, 2003) to create an initial network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$. Connections within each cluster are formed with a probability of $0.5$, while inter-cluster connections occur with a probability of $0.1$. This setup mirrors our assumption that nodes within the same cluster are more inclined to connect due to a higher number of shared neighbors. In subsequent time steps, we then examine each node $i$, considering the intersection of neighborhoods of $i$’s non-neighbors or the number of mutual friends between $i$ and any non-neighbor $j$ [^5] It’s worth noting that similar outcomes arise when providing neighbors instead of common neighbors.
We conducted ten simulations with $n=50$ nodes to facilitate clear visualization and ensure statistical significance [^6] Choosing $n=50$ instead of a larger number like $n=200$ aids in visualization and maintains statistical significance.. To contrast LLM-based network formation, we refer to a random null model, selecting nodes at random for connection.
For evaluating triadic closure, we utilize two metrics: marginal transitivity and probability of edge formation within the same community. Marginal transitivity ($D$) represents the change in the ratio of closed triangles to all triads, transitioning from the initial network $G_{T}$ to the SBM-generated $G_{1}$:
where a large positive $D$ indicates a strong triadic closure tendency.
The probability of forming an edge within the same community ($\hat{p}$) is calculated by the ratio of edges in $G_{T}\setminus G_{1}$ (newly formed edges) connecting nodes within the same cluster:
@@
where $y_{i},y_{j}\in{A,B}$ denote the community memberships of nodes $i$ and $j$, respectively. A value of $\hat{p}$ exceeding $0.5$ suggests a triadic closure tendency. As we investigate under SBM, same community membership indicates more open triads being closed.
In Figure 2(a), sample networks are displayed, with the upper panel showing networks where the entire structure is provided and the lower panel showing those with only common neighbor numbers provided. Nodes are color-coded to indicate their cluster memberships in the SBM, with red and blue edges within clusters and orange edges between clusters. Newly formed edges are highlighted with thicker lines.
Marginal transitivity, presented in the fourth panel of each row, shows a statistically significant increase across all temperatures ($P<0.001$, t-test), and compared to the Random model, the increase is notable ($P<0.001$, t-test).
Regarding the probability of within-community network formation (last panels in each row), we found that across all temperatures, this probability significantly exceeds that in the random network formation scenario ($P<0.001$, t-test against Random), indicating an LLM tendency towards triadic closure.
These findings affirm that LLM agents tend to follow triadic closure. We investigate the reasons provided by LLMs for their decisions in SI Section 3.777Reasons for decisions based on the number of common neighbors are not provided, as these decisions are solely based on this information. The analysis revealed that many decisions were influenced by the number of mutual friends.
@@
To measure the presence of homophily, we measure the attribute assortativity coefficient of each of the three features (hobby, color, location). For each property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$ (indexed by $k$ or $l$), its assortativity coefficient $R$ is defined as
$R=\frac{\sum_{k=1}^{K}M_{kk}-\sum_{k=1}^{K}a_{k}b_{k}}{1-\sum_{k=1}^{K}a_{k}b_{k}}.$
Here $M$ represents the mixing matrix. Its elements $M_{kl}$ reflect the proportion of edges connecting two nodes with values $P_{k}$ and $P_{l}$, respectively. We define $a_{k}=\sum_{l=1}^{K}M_{kl}$ and $b_{k}=\sum_{l=1}^{K}M_{lk}$. Assortativity ranges from $-1$ to $+1$. A positive assortativity indicates nodes preferentially connect to similar ones, forming a homophilous network. Conversely, a negative assortativity suggests connections primarily occur between dissimilar nodes, indicating heterophily.
In Figure 3(a), we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures. Moreover, the observed assortativity at all temperatures significantly differs from Random, which has assortativity close to 0 ($P<0.001$, t-test).
Moreover, we present the probability of each reason being mentioned in SI Section 3. We find that shared hobbies appear to be much more important than shared location or favorite color. This is consistent with our intuition that, similarly to humans, LLMs care more about hobbies than other superficial attributes when forming social connections. In token, we conclude homophily emerges for the networks generated by LLMs.
Macro-Level Principles
(a) Assortativity Coefficients
(b) Network Instances
(c) Communities Statistics
Figure 3. Results for Principle 3 (homophily). (a) Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily. The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random). (b) Networks created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. The edge widths correspond to the number of common attributes shared between the nodes. The node colors correspond to the communities found by running the Louvain algorithm. (c) Louvain modularity, number of communities, and community sizes of networks created based on Principle 3. Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test). For smaller temperatures, the communities are more “well-separated”. The error bars correspond to 95% confidence intervals.
@@
Both triadic closure and homophily contribute to the formation of community structures. By examining how these two factors contribute to network formation, we aim to gain insights into the underlying mechanisms driving community dynamics in LLM-generated networks. We employ the simulation results presented in the synthetic networks to determine whether community structure in networks generated by LLMs emerges from triadic closure or homophily.